# Approximating Discrimination Within Models When Faced With Several Non-Binary Sensitive Attributes

### v1

@@ 4.1 Experimental setups
Implementation details
We mainly use bagging, AdaBoost, LightGBM [40], FairGBM [39], and AdaFair [41] as learning algorithms, where FairGBM and AdaFair are two fairness-aware ensemble-based methods. Plus, certain kinds of classifiers are used in Section 4.2—including decision trees (DT), naive Bayesian (NB) classifiers, $k$-nearest neighbours (KNN) classifiers, Logistic Regression (LR), support vector machines (SVM), linear SVMs (linSVM), and multilayer perceptrons (MLP)—so that we have a larger learner pools to choose from based on different fairness-relevant rules. Standard 5-fold cross-validation is used in these experiments, in other words, in each iteration, the entire dataset is divided into two parts, with 80% as the training set and 20% as the test set. Also, features of datasets are scaled in preprocessing to lie between 0 and 1. Except for the experiments for RQ3, we set the hyper-parameters $m_{1}=25$ and $m_{2}=\lceil 2\lg(n)\rceil$ in other experiments.
4.2 Comparison between HFM and baseline fairness measures
The aim of this experiment is to evaluate the effectiveness of the proposed HFM compared with baseline fairness measures. As groundtruth discriminative levels of classifiers remain unknown and it is hard to directly compare different methods from that perspective, we compare the correlation (referring to the Pearson correlation coefficient) between the performance difference and different fairness measures. The empirical results are reported in Figures 12 and Tables IIIIII.
For one single sensitive attribute, we can see from Figure 1 that $\mathbf{df}^{\text{avg}}$ is highly correlated with recall/sensitivity and $\mathrm{f}_{1}$ score. Besides, even $\mathbf{df}^{\text{avg}}$ only describes the extra bias, its correlation with $\Delta($performance)))) is still close to that of DR (and sometimes DP), which means HFM can capture the bias within classifiers indeed and that HFM captures it more finely than our previous work [34]. Moreover, $\mathbf{df}^{\text{avg}}$ shows higher correlation with $\Delta($performance)))) than $\mathbf{df}$ in most cases, which means $\mathbf{df}^{\text{avg}}$ may capture the extra bias level of classifiers better than $\mathbf{df}$ in practice.
As for multiple sensitive attributes, we can see from Figure 2 that $\mathbf{df}^{\text{avg}}$ is highly correlated with recall/sensitivity and $\mathrm{f}_{1}$ score and that $\mathbf{df}^{\text{avg}}$ shows higher correlation with $\Delta($performance)))) than $\mathbf{df}$ in most cases, which is similar to our observation in Figure 1. Note that the original DR [32] is only for one single sensitive attributes with binary or multiple values, and for comparison with HFM, we calculate $\text{DR}_{\text{avg}}=\frac{1}{n_{a}}\sum_{i=1}^{n_{a}}\text{DR}_{i}$ here, analogously to $\mathbf{df}^{\text{avg}}$. Besides, we observe that the correlation between $\mathbf{df}^{\text{avg}}$ and $\Delta\text{Accuracy}$ (resp. $\Delta\mathrm{f}_{1}$ score, $\Delta\text{Specificity}$) achieves half of that of DR, and $\mathbf{df}^{\text{avg}}$ even outperforms DR concerning $\Delta\text{Recall}$. Given that HFM only captures the extra bias introduced by classifiers, we believe at least $\mathbf{df}^{\text{avg}}$ could capture quite a part of bias within.
Furthermore, we report plots of fairness-performance trade-offs per fairness measure in Fig. 4. We can see that: 1) for one single sensitive attribute, HFM (i.e., $\mathbf{df}$ and $\mathbf{df}^{\text{avg}}$) achieves the best result in Fig. 4LABEL:sub@subfig:tradeoff,a and 4LABEL:sub@subfig:tradeoff,c; and 2) for all sensitive attributes on one dataset, $\mathbf{df}$ and $\mathbf{df}^{\text{avg}}$ perform closely and both outperform $\text{DR}_{\text{avg}}$ in Fig. 4LABEL:sub@subfig:tradeoff,b and 4LABEL:sub@subfig:tradeoff,d. This observation demonstrates the effectiveness of HFM from another perspective, in other words, HFM could work well if fairness-performance trade-offs need to be considered.
In this subsection, we evaluate the performance of the proposed ApproxDist and ExtendDist compared with the precise distance that is directly calculated by definitions. To verify whether they could achieve the true distance between sets precisely and timely, we employ scatter plots to compare their values and time cost, presented in Fig. 4. Note that $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ and $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ are computed together in ApproxDist at one time, and so are $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ and $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ in ExtendDist. Also notice that the previous ApproxDist [34] is included for comparison to its current version in scenarios of binary values.
4.3.1 Validity of ApproxDist
As we can see from Figures 4LABEL:sub@subfig:approx,3 and 4LABEL:sub@subfig:approx,4, the approximated values of maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ are highly correlated with their corresponding precise values. Besides, their linear fit line and the identity line (that is, $f(x)=x$) are near and almost parallel, which means the approximated values are pretty close to their precise value. Similar observations are concluded for the average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ shown in Figures 4LABEL:sub@subfig:approx,7 and 4LABEL:sub@subfig:approx,8. As for the execution time of approximation and direct computation in Figures 4LABEL:sub@subfig:approx,11 and 4LABEL:sub@subfig:approx,12, ApproxDist may take a bit longer time in scenarios of multi-value cases than that of binary values, while all of them could achieve a shorter time than precise values when the execution of direct computation is costly.
4.3.2 Validity of ExtendDist
As we can see from Figures 4LABEL:sub@subfig:approx,1 and 4LABEL:sub@subfig:approx,2, the approximated values of maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ are highly correlated with their corresponding precise values. Besides, their linear fit line and the identity line are near and almost parallel, which means the approximated values are pretty close to their precise value. Similar observations are concluded for the average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S)$ shown in Figures 4LABEL:sub@subfig:approx,5 and 4LABEL:sub@subfig:approx,6. As for the execution time of approximation and direct computation in Figures 4LABEL:sub@subfig:approx,9 and 4LABEL:sub@subfig:approx,10, ExtendDist would obtain a bigger advantage when computing precise values is expensive, while on the opposite, we do not need ExtendDist that much and can directly calculate them instead.
4.4 Effect of hyper-parameters $m_{1}$ and $m_{2}$
@@
In this subsection, we investigate whether different choices of hyper-parameters (that is, $m_{1}$ and $m_{2}$) would affect the performance of ApproxDist and ExtendDist or not. Different $m_{2}$ values are tested when $m_{1}$ is fixed, and vice versa, with empirical results presented in Figures 6 and 6.
@@ 4.4.1 Effect on ApproxDist
As we can see from Figures 6LABEL:sub@subfig:pm,ext,9 and 6LABEL:sub@subfig:pm,ext,11, when direct computation of distances (i.e., maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ and average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$) is expensive, obtaining their approximated values via ApproxDist distinctly costs less time than that of precise values by Eq. (4) and (5). Increasing $m_{2}$ (or $m_{1}$) in ApproxDist would cost more time, while the effect of increasing $m_{1}$ is more obvious.
As for the approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ shown in Fig. 6LABEL:sub@subfig:pm,ext,1 and 6LABEL:sub@subfig:pm,ext,3 as well as approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ shown in Fig. 6LABEL:sub@subfig:pm,ext,5 and 6LABEL:sub@subfig:pm,ext,7, all approximated values are highly correlated and close to the precise values of distance no matter how small $m_{2}$ (or $m_{1}$) is, which means the effect of improper choices of hyper-parameters is unapparent; As $m_{2}$ increases, the approximated values would be closer to the precise values of distance, while the effect of changing $m_{1}$ would be less manifest.
@@ 4.4.2 Effect on ExtendDist
As we can see from Figures 6LABEL:sub@subfig:pm,ext,10 and 6LABEL:sub@subfig:pm,ext,12, when direct computation of distances (i.e., maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ and average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S)$) is expensive, obtaining their approximated values via ExtendDist distinctly costs less time than that of precise values by Eq. (6) and (7). Increasing $m_{2}$ (or $m_{1}$) in ExtendDist would cost more time, while the effect of increasing $m_{1}$ is more obvious.
As for the approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ shown in Figures 6LABEL:sub@subfig:pm,ext,2 and 6LABEL:sub@subfig:pm,ext,4 as well as approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S)$ shown in Figures 6LABEL:sub@subfig:pm,ext,6 and 6LABEL:sub@subfig:pm,ext,8, all approximated values are highly correlated and close to the precise values of distance no matter how small $m_{2}$ (or $m_{1}$) is, which means the effect of improper choices of hyper-parameters is unapparent; As $m_{2}$ increases, the approximated values would be closer to the precise values of distance, while the effect of changing $m_{1}$ would be less manifest.
@@ 4.4.3 Comparison of ApproxDist between our previous work [34] and the current version in this paper
We can see from Figures 6LABEL:sub@subfig:pm,bin,1, 6LABEL:sub@subfig:pm,bin,2, 6LABEL:sub@subfig:pm,bin,5, and 6LABEL:sub@subfig:pm,bin,6 that our previous ApproxDist demonstrates slightly higher correlation to precise values of maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ (aka. $\mathbf{D}(S_{1},\bar{S}_{1})$ in scenarios of binary values) than its current version in this work; Different choices of $m_{2}$ (or $m_{1}$) cause nearly imperceptible effect on their approximation effectiveness. The previous version also shows close and even better performance on compressed time cost than the current ApproxDist here, depicted in Figures 6LABEL:sub@subfig:pm,bin,3, 6LABEL:sub@subfig:pm,bin,4, 6LABEL:sub@subfig:pm,bin,7, and 6LABEL:sub@subfig:pm,bin,8, especially when direct computation is not much expensive. However, when the execution time cost of direct computation is relatively cheap, ExtendDist displays messy and worse execution speed than ApproxDist, shown in Figures 6LABEL:sub@subfig:pm,ext,11 and 6LABEL:sub@subfig:pm,ext,12. We believe there are mainly two reasons for this phenomenon: one is that AcceleDist is repeated twice from line 2 to line 5 in Algorithm 2 while it is executed only once in the previous ApproxDist [34], causing Algorithm 2 a slightly longer execution time than its previous version; the other is that parallel computing is integrated in ExtendDist in practice to further accelerate its execution, detailed more in Section 4.5 and Figure 7.
@@ 4.5 Discussion and limitations
Given the wide applications of ML models in the real world nowadays and the complexity of discrimination mitigation in the face of multiple factors interweaving, it matters a lot to bring in such techniques to deal with several sensitive attributes with even multiple values. Therefore, our work provides a fine-grained fairness measure option named HFM that captures the bias level of models more finely, in order to better detect and moderate discrimination within. The proposed HFM are suitable for both binary and multi-class classification, thus enlarging its applicable value. To promptly approximate the value of HFM, we further proposed ApproxDist and ExtendDist to speed up the expensive calculation process, of which the effectiveness and efficiency have been demonstrated in Section 4.3. However, there are also limitations in the proposed approximation algorithms. The major one is that their time cost will significantly increase if the number of optional values within a sensitive attribute is relatively large. For instance, the computation incurring on the PPR/PPVR datasets may take close or sometimes even longer time than that on the Income dataset, even though the latter has way more instances than the former, because there are six sub-groups under the race attribute on PPR/PPVR while the number is only five on Income. Therefore, we integrate parallel computing in practice to further raise the execution speed of ExtendDist. To be specific, we use three cores to run lines 1 to 3 of Algorithm 1 in parallel in our experiments (Fig. 7), while the choice of the number of cores is not a fixed constant. In other words, using two or four cores is also acceptable if the practitioners like. Furthermore, it is easy to tell that there is still room for improvement in the approximation algorithms. For instance, it might achieve a shorter time of computing in ApproxDist if the procedure between line 1 and line 8 in Algorithm 2 could be executed in parallel, although we did not perform that this time. So does that between line 3 and line 11 of Algorithm 3. But we believe it will need a more deliberate design to balance the parallel computing among them in case the cost of generating more threads/processes to achieve it is not worthy as expected compared with its computing results. Therefore, we would rather leave it in future work, instead of cramming too much and confusing our main contributions in this work.
@@ 5 Conclusion
In this paper, we investigate how to evaluate the discrimination level of classifiers in the face of multi-attribute protection scenarios and present a novel harmonic fairness measure with two optional versions (that is, maximum HFM and average HFM), of which both are based on distances between sets from a manifold perspective. To accelerate the computation of distances between sets and reduce its time cost from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$, we further propose two approximation algorithms (that is, ApproxDist and ExtendDist) to resolve bias evaluation in scenarios for single attribute protection and multi-attribute protection, respectively. Furthermore, we provide an algorithmic effectiveness analysis for ApproxDist under certain assumptions to explain how well it could work theoretically. The empirical results have demonstrated that the proposed fairness measure (including maximum HFM and average HFM) and approximation algorithms (i.e., ApproxDist and ExtendDist) are valid and effective.

### v2

@@ 4.3.3 Comparison with EarlyBreak [48]
Figure 9: Comparison between ApproxDist and EarlyBreak [48]. Note that ApproxDist (prev) is the approximation algorithm from our previous work [36], which is also supposed to approximate $\mathbf{D}_{\cdot}(S_{1},\bar{S}_{1})$, equivalent to $\mathbf{D}_{\cdot,\mathop{\bm{a}}}(S,a_{i})$ in this paper.
4.4 Effect of hyperparameters $m_{1}$ and $m_{2}$
@@
In this subsection, we investigate whether different choices of hyperparameters (that is, $m_{1}$ and $m_{2}$) would affect the performance of ApproxDist and ExtendDist or not. Different $m_{2}$ values are tested when $m_{1}$ is fixed, and vice versa, with empirical results presented in Figures 7 and 7.
@@ 4.4.1 Effect on ApproxDist
As we can see from Figures 7LABEL:sub@subfig:pm,ext,9 and 7LABEL:sub@subfig:pm,ext,11, when direct computation of distances (i.e. maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ and average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$) is expensive, obtaining their approximated values via ApproxDist distinctly costs less time than that of precise values by Eq. (4) and (5). Increasing $m_{2}$ (or $m_{1}$) in ApproxDist would cost more time, while the effect of increasing $m_{1}$ is more obvious.
As for the approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ shown in Fig. 7LABEL:sub@subfig:pm,ext,1 and 7LABEL:sub@subfig:pm,ext,3 as well as approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ shown in Fig. 7LABEL:sub@subfig:pm,ext,5 and 7LABEL:sub@subfig:pm,ext,7, all approximated values are highly correlated and close to the precise values of distance, no matter how small $m_{2}$ (or $m_{1}$) is, which means the effect of improper choices of hyperparameters is unapparent; As $m_{2}$ increases, the approximated values would be closer to the precise values of distance, while the effect of changing $m_{1}$ would be less manifest.
@@ 4.4.2 Effect on ExtendDist
As we can see from Figures 7LABEL:sub@subfig:pm,ext,10 and 7LABEL:sub@subfig:pm,ext,12, when direct computation of distances (i.e. maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ and average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S)$) is expensive, obtaining their approximated values via ExtendDist distinctly costs less time than that of precise values by Eq. (6) and (7). Increasing $m_{2}$ (or $m_{1}$) in ExtendDist would cost more time, while the effect of increasing $m_{1}$ is more obvious.
As for the approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ shown in Figures 7LABEL:sub@subfig:pm,ext,2 and 7LABEL:sub@subfig:pm,ext,4 as well as approximation performance of $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S)$ shown in Figures 7LABEL:sub@subfig:pm,ext,6 and 7LABEL:sub@subfig:pm,ext,8, all approximated values are highly correlated and close to the precise values of distance no matter how small $m_{2}$ (or $m_{1}$) is, which means the effect of improper choices of hyperparameters is unapparent; As $m_{2}$ increases, the approximated values would be closer to the precise values of distance, while the effect of changing $m_{1}$ would be less manifest.
@@ 4.4.3 Comparison of ApproxDist between our previous work [36] and the current version in this paper
We can see from Figures 7LABEL:sub@subfig:pm,bin,1, 7LABEL:sub@subfig:pm,bin,2, 7LABEL:sub@subfig:pm,bin,5, and 7LABEL:sub@subfig:pm,bin,6 that our previous ApproxDist demonstrates slightly higher correlation to precise values of maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ (also known as $\mathbf{D}(S_{1},\bar{S}_{1})$ in scenarios of binary values) than its current version in this work; Different choices of $m_{2}$ (or $m_{1}$) cause nearly imperceptible effects on their approximation effectiveness. The previous version also shows close and even better performance on compressed time cost than the current ApproxDist here, depicted in Figures 7LABEL:sub@subfig:pm,bin,3, 7LABEL:sub@subfig:pm,bin,4, 7LABEL:sub@subfig:pm,bin,7, and 7LABEL:sub@subfig:pm,bin,8, especially when direct computation is not much expensive. However, when the execution time cost of direct computation is relatively cheap, ExtendDist displays a messy and worse execution speed than ApproxDist, shown in Figures 7LABEL:sub@subfig:pm,ext,11 and 7LABEL:sub@subfig:pm,ext,12. We believe there are mainly two reasons for this phenomenon: one is that AcceleDist is repeated twice from line 2 to line 5 in Algorithm 2 while it is executed only once in the previous ApproxDist [36], causing Algorithm 2 a slightly longer execution time than its previous version; the other is that parallel computing is integrated in ExtendDist in practice to further accelerate its execution, detailed more in Section 4.6 and Fig. 8.
Figure 10: Potential exploration for more complex features, using a multilayer perceptron (MLP). Note that $\mathbf{df}_{\text{prev}},\mathbf{df},\mathbf{df}^{\text{avg}}$ are calculated based on original features, while $\mathbf{df}_{\text{prev,emb}},\mathbf{df}_{\text{emb}},\mathbf{df}_{\text{emb}}^{\text{avg}}$ are calculated using replaced features in the MLP, as described in Section 4.5. Also note that $\hat{\mathbf{df}}$ in (c) and (d) indicates that the distances are calculated using approximation algorithms, while $\mathbf{df}$ indicates the distances are calculated using the direct computation; So are the others.
4.5 Potential exploration for more complex feature sets
We did some preliminary exploration into the possibility of applying HFM to more complex feature sets, such as visual features. Instead of directly using HFM on visual datasets, we use the represented (or embedded) features before the final output of a neural network as an alternative way to test it. In other words, we still compute $\mathbf{D}{}_{\mathop{\bm{a}}}(S,a_{i})$ with the original features, but replace those features with the represented vectors in neural networks in $\mathbf{D}{}_{f,\mathop{\bm{a}}}(S,a_{i})$, and then calculate $\mathbf{D}{}_{\mathop{\bm{a}}}(S)$, $\mathbf{D}{}_{f,\mathop{\bm{a}}}(S)$, and $\mathbf{df}(f)$ as usual. Analogously, we can get $\mathbf{D}^{\text{avg}}_{f,\mathop{\bm{a}}}(S,a_{i})$, $\mathbf{D}^{\text{avg}}_{f,\mathop{\bm{a}}}(S)$, and $\mathbf{df}^{\text{avg}}(f)$. The neural network we used in this experiment was a multilayer perceptron, where the input dimension depends on the dataset in use, followed by one hidden layer with a size of 256, ReLU, one hidden layer with a size of 64, ReLU, and one fully connected layer before Softmax. The empirical results are presented in Fig. 10, where we can observe that $\mathbf{df}_{\text{emb}}$ using the embedded features shows a higher correlation with $\Delta$Accuracy, compared with $\mathbf{df}$, meaning that HFM has the potential to be applied to complex features.
@@ 4.6 Discussion and limitations
Given the wide applications of ML models in the real world nowadays and the complexity of discrimination mitigation in the face of multiple factors interweaving, it matters a lot to bring in such techniques to deal with several SAs with even multiple values. Therefore, our work provides a fine-grained fairness measure option named HFM that captures the bias level of models more finely, in order to better detect and moderate discrimination within. The proposed HFM is suitable for both binary and multi-class classification, thus enlarging its applicable value. To efficiently approximate the value of HFM, we further proposed ApproxDist and ExtendDist to speed up the expensive calculation process, of which the effectiveness and efficiency have been demonstrated in Section 4.3. However, there are also limitations in the proposed approximation algorithms. The major one is that their time cost will significantly increase if the number of possible values within a SA is relatively large. For instance, the computation incurring on the PPR/PPVR datasets may take close or sometimes even longer time than that on the Income dataset, even though the latter has way more instances than the former, because there are six sub-groups under the race attribute on PPR/PPVR while the number is only five on Income. Therefore, we integrate parallel computing in practice to further raise the execution speed of ExtendDist. To be specific, we use three cores to run lines 1 to 3 of Algorithm 1 in parallel in our experiments (Fig. 8), while the choice of the number of cores is not a fixed constant. In other words, using two or four cores is also acceptable if the practitioners like it. Furthermore, it is easy to tell that there is still room for improvement in the approximation algorithms. For instance, it might achieve faster computing times in ApproxDist if the procedure between lines 1 and 8 in Algorithm 2 could be executed in parallel, although we did not perform that this time. So does that between lines 3 and 11 of Algorithm 3. But we believe it will need a more deliberate design to balance the parallel computing among them in case the cost of generating more threads/processes to achieve it is not worth enough as expected compared with its computing results. Therefore, we would rather leave it for future work, instead of cramming too much and confusing our main contributions in this work.
@@ 5 Conclusion
In this paper, we investigate how to evaluate the discrimination level of classifiers in the face of multi-attribute protection scenarios and present a novel harmonic fairness measure with two optional versions (that is, maximum HFM and average HFM), of which both are based on distances between sets from a manifold perspective. To accelerate the computation of distances between sets and reduce its time cost from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$, we further propose two approximation algorithms (that is, ApproxDist and ExtendDist) to resolve bias evaluation in scenarios for single attribute protection and multi-attribute protection, respectively. Furthermore, we provide an algorithmic effectiveness analysis for ApproxDist under certain assumptions to explain how well it could work theoretically. The empirical results have demonstrated that the proposed fairness measure (including maximum HFM and average HFM) and approximation algorithms (i.e. ApproxDist and ExtendDist) are valid and effective.